# Integrating representative and non-representative survey data for efficient inference

### v1

@@ 3 Simulation study
To assess the extent to which the synthesis method reduces the now-casted RMSE and the effects of mis-specifying the generating mechanism of $\phi_{kt}$, a 3 x 3 simulation design is performed. We consider 3 generating mechanisms for $\phi_{kt}$, where $\phi_{kt}$ either is constant in time (7), linear in time (8), or follows a random walk (9), all as described in the methods section. To get a good representation of how the methods perform under many parameter conditions we generate the parameters from the prior distributions. That is, each simulated dataset arrises from different underlying parameter values. One issue from this technique is that priors presented in the methods section are wider than what might be realistic. For example, if one or more of the parameters are extreme in value, then $Y_{kt}$ may be equal to 0 or $n_{kt}$ for many values of $t$. To avoid this, we narrow the priors to more plausible values:
@@
This translates to $\textrm{logit}^{-1}(\theta_{0})$ having roughly 90% of its prior density between 0.22 and 0.78. Meaning most of the time we believe the initial positive rate is near 50%. We also set narrow our belief about $\sigma^{2}$ which means $\textrm{logit}^{-1}(\theta_{t})-\textrm{logit}^{-1}(\theta_{t-1})$ is $\pm$0.15 or less, 90% of the time, assuming $\textrm{logit}^{-1}(\theta_{t-1})=0.5$. For the linear model, $\phi_{kt}-\phi_{k(t-1)}$ in between -0.33 and 0.33 90% of the time when $\phi_{k(t-1)}=1$ Lastly, for the random walk model $\phi_{kt}-\phi_{k(t-1)}$ is between -0.37 and 0.40, 90% of the time, where $\phi_{k(t-1)}=1$. We simulate the data with $K=3$ surveys and with a population of $N=10,000,000$. Let $k=1$ be the unbiased survey, and $k=2$ and $3$ be the non-probability surveys. Given that non-probability surveys are often larger we set $n_{1t}=100$, and $n_{2t}=n_{3t}=1000$ for all $t$. We repeat this experiment for 3 sets of time-points: ($T$ = 5, 10, and 15). The models are all fit in JAGS 4.3.x (Plummer et al. 2003), with roughly 20,000-25,000 burn-in and a sample size of 50,000-70,000 depending on the complexity of $\phi_{kt}$. The goal is to compare the RMSE, i.e., the root averaged squared difference between $\textrm{logit}^{-1}(\theta_{T})$ and the posterior median across the 1000 parameters generated.
@@ 3.2 Simulation Results
Figure 2: Results of simulation study for t = 5, 10, and 15 time-points. Simulations run using the model fit in JAGS 4.3.x. The RMSE is calculated as the root estimated averaged difference between $\textrm{logit}^{-1}(\theta_{T})$ and the posterior median.
Figure 2 shows the results of the simulation study. As expected, when the model for $\phi_{kt}$ is correctly specified, the RMSE is the lowest. Moreover, incorporating the additional 2 biased surveys generally leads to a lower RMSE, showing the information from the biased surveys is useful. However, if the model for $\phi_{kt}$ is drastically mis-specified, the synthesis model performs worse than using the unbiased survey in isolation. Specifically, fitting a constant $\phi_{kt}$, when $\phi_{kt}$ actually changes significantly, leads to a large increase in RMSE. These simulations also suggest that the added complexity of fitting a random walk $\phi_{kt}$ does not lead to a large increase in estimator variance. Surprisingly, for scenarios with even a few time-points, a random walk model for $\phi_{kt}$ appears to be a relatively safe assumption. By increasing the number of time-points, additional complexity can be used to better adjust for the bias without gaining excessive estimator variance.
@@ 4.4 Inclusion of biased surveys quantified in terms of increased unbiased sample size
Then we can easily get the number of iid samples required for a $(1-\alpha/2)100\%$ confidence interval:
$n_{\text{iid samples }t}=\frac{Z_{\alpha/2}^{2}\hat{p}_{t}(1-\hat{p}_{t})}{R_{t}\text{MOE}_{t}}.$ (7)
Figure 8: Plot of number of iid samples gained by date, for each survey. The bars represent $n_{\text{iid Synthesis}}-n_{\text{iid Axios-Ipsos}}$ at the 95% confidence level.
@@
Table 2: Summary statistics on the effective number of gained iid samples from the inclusion of the biased surveys, rounded to the nearest integer.
Figure 8 shows the plots of the gain in iid samples by time-point. We can see that there is considerable information gained for later time-points. The first and fourth time-point have small negative gains when the Household-Pulse survey is included. This is likely due to there being two missing values in four time-points for the Household-Pulse survey (and one missing Axios-Ipsos time-point in between) leading to large posterior variances for parameters. Thus many of these bars are not completely accurate given many of the surveys are missing for some time-points, and we would expect the gain to be larger if all three surveys were present for all time-points. Ultimately, synthesizing a larger survey (in terms of sample size) translates into a larger gain in $n_{\textrm{iid}}$ for the synthesis method. Meaning we could conceptually achieve the properties of an unbiased large sample size survey by combing a small unbiased survey with an extremely large biased survey.
@@ 5.2 Selecting a model for $\phi$
In this paper we have provided three possible models for how $\phi$ may evolve over time: the constant model, the linear model, and the random walk model. The choice of selecting the model for $\phi$ depends on the prior belief about the evolution of the bias and the bias variance trade-off of mis-specifying the model. In most applications the bias would neither be constant nor linear, but interpolating the crude error of a random walk with a linear model for $\phi$ when there are few time-points is a reasonable choice. However, as shown in Figure 2, choosing a random walk model is a conservative and surprisingly efficient choice. There is also the possibility of extending the complexity of the model for $\phi$ if the number of time-points is large enough. We could consider employing a moving average model or adding other higher order auto-regressive terms. Yet, we must also be careful about adding too much complexity to a latent parameter, but the possibility exists. Furthermore, we have only considered fitting the same model for both of the non-representative surveys. There is an argument to be made to either specify different models for $\phi_{k}$ for each $k$, or to force the non-representative surveys to share the same jump or slope. In the latter case, this assumes the bias $\phi$ is not necessarily a parameter of the survey, but a parameter of the population that surveys sample from. In either case, the most conservative model (random walk) shown in this paper can be relied upon, having been shown to have a lower RMSE than using the unbiased survey, even when mis-specified.

### v2

@@ 3 Simulation study
To assess the extent to which the synthesis method reduces the now-casted MSE and the effects of mis-specifying the generating mechanism of $\phi_{kt}$, a 3 x 3 simulation design is performed. We consider 3 generating mechanisms for $\phi_{kt}$, where $\phi_{kt}$ either is constant in time (7), linear in time (8), or follows a random walk (9), all as described in the methods section. To get a good representation of how the methods perform under many parameter conditions we generate the parameters from the prior distributions. That is, each simulated dataset arises from different underlying parameter values. One issue from this technique is that priors presented in the methods section are wider than what might be realistic. For example, if one or more of the parameters are extreme in value, then $Y_{kt}$ may be equal to 0 or $n_{kt}$ for many values of $t$. To avoid this, we narrow the priors to more plausible values:
@@
This translates to $\textrm{logit}^{-1}(\theta_{0})$ having roughly 90% of its prior density between 0.22 and 0.78. Meaning most of the time we believe the initial positive rate is near 50%. We also set narrow our belief about $\sigma^{2}$ which means $\textrm{logit}^{-1}(\theta_{t})-\textrm{logit}^{-1}(\theta_{t-1})$ is $\pm$0.15 or less, 90% of the time, assuming $\textrm{logit}^{-1}(\theta_{t-1})=0.5$. For the linear model, $\phi_{kt}-\phi_{k(t-1)}$ in between -0.33 and 0.33 90% of the time when $\phi_{k(t-1)}=1$ Lastly, for the random walk model $\phi_{kt}-\phi_{k(t-1)}$ is between -0.37 and 0.40, 90% of the time, where $\phi_{k(t-1)}=1$. We simulate the data with $K=3$ surveys and with a population of $N=10,000,000$. Let $k=1$ be the unbiased survey, and $k=2$ and $3$ be the non-probability surveys. Given that non-probability surveys are often larger we set $n_{1t}=100$, and $n_{2t}=n_{3t}=1000$ for all $t$. We repeat this experiment for 3 sets of time-points: ($T$ = 5, 10, and 15). The models are all fit in JAGS 4.3.x (Plummer et al. 2003), with roughly 20,000-25,000 burn-in and a sample size of 50,000-70,000 depending on the complexity of $\phi_{kt}$. The goal is to compare the MSE, i.e., the root averaged squared difference between $\textrm{logit}^{-1}(\theta_{T})$ and the posterior median across the 2000 parameters generated. We add 95% confidence intervals for the Monte Carlo MSE estimates based on the central limit theorem and the MCSE (Monte Carlo Standard Error), see Morris et al. (2019).
@@ 3.2 Simulation Results
Figure 2: Results of simulation study for t = 5, 10, and 15 time-points. Simulations run using the model fit in JAGS 4.3.x. The MSE is calculated as the estimated averaged squared difference between $\textrm{logit}^{-1}(\theta_{T})$ and the posterior median. Error bars represent the 95% confidence intervals for the Monte Carlo estimated MSE i.e. $\hat{\textrm{MSE}}\pm 1.96\textrm{MCSE}$.
Figure 2 shows the results of the simulation study. As expected, when the model for $\phi_{kt}$ is correctly specified, the MSE is the lowest. Moreover, incorporating the additional 2 biased surveys generally leads to a lower MSE, showing the information from the biased surveys is useful. However, if the model for $\phi_{kt}$ is drastically mis-specified, the synthesis model performs worse than using the unbiased survey in isolation. Specifically, fitting a constant $\phi_{kt}$, when $\phi_{kt}$ actually changes significantly, leads to a large increase in MSE. These simulations also suggest that the added complexity of fitting a random walk $\phi_{kt}$ does not lead to a large increase in estimator variance. Surprisingly, for scenarios with even a few time-points, a random walk model for $\phi_{kt}$ appears to be a relatively safe assumption. By increasing the number of time-points, additional complexity can be used to better adjust for the bias without gaining excessive estimator variance. Moreover, the confidence intervals of for Monte Carlo estimated MSE are sufficiently separated to have confidence in the results.
@@ 4.4 Inclusion of biased surveys quantified in terms of increased unbiased sample size
Then we can easily get the number of $n_{\textrm{iid}}$ required for a $(1-\alpha/2)100\%$ confidence interval:
$n_{\text{iid}t}=\frac{Z_{1-\alpha/2}^{2}\hat{p}_{t}(1-\hat{p}_{t})}{R_{t}\text{MOE}_{t}}.$ (7)
Figure 8: Plot of number of the number of $n_{\textrm{iid}}$ gained by date, for each survey. The bars represent $n_{\text{iid Synthesis}}-n_{\text{iid Axios-Ipsos}}$ at the 95% confidence level.
@@
Table 2: Summary statistics on the effective number of $n_{\textrm{iid}}$ gained from the inclusion of the biased surveys, rounded to the nearest integer.
Figure 8 shows the plots of the gain in $n_{\textrm{iid}}$ by time-point. We can see that there is considerable information gained for later time-points. The first and fourth time-point have small negative gains when the Household-Pulse survey is included. This is likely due to there being two missing values in four time-points for the Household-Pulse survey (and one missing Axios-Ipsos time-point in between) leading to large posterior variances for parameters. Thus many of these bars are not completely accurate given many of the surveys are missing for some time-points, and we would expect the gain to be larger if all three surveys were present for all time-points. Ultimately, synthesizing a larger survey (in terms of sample size) translates into a larger gain in $n_{\textrm{iid}}$ for the synthesis method. Meaning we could conceptually achieve the properties of an unbiased large sample size survey by combing a small unbiased survey with an extremely large biased survey.
@@ 5.2 Selecting a model for $\phi$
In this paper we have provided three possible models for how $\phi$ may evolve over time: the constant model, the linear model, and the random walk model. The choice of selecting the model for $\phi$ depends on the prior belief about the evolution of the bias and the bias variance trade-off of mis-specifying the model. In most applications the bias would neither be constant nor linear, but interpolating the crude error of a random walk with a linear model for $\phi$ when there are few time-points is a reasonable choice. However, as shown in Figure 2, choosing a random walk model is a conservative and surprisingly efficient choice. There is also the possibility of extending the complexity of the model for $\phi$ if the number of time-points is large enough. We could consider employing a moving average model or adding other higher order auto-regressive terms. Yet, we must also be careful about adding too much complexity to a latent parameter, but the possibility exists. Furthermore, we have only considered fitting the same model for both of the non-representative surveys. There is an argument to be made to either specify different models for $\phi_{k}$ for each $k$, or to force the non-representative surveys to share the same jump or slope. In the latter case, this assumes the bias $\phi$ is not necessarily a parameter of the survey, but a parameter of the population that surveys sample from. In either case, the most conservative model (random walk) shown in this paper can be relied upon, having been shown to have a lower MSE than using the unbiased survey, even when mis-specified.